OpenStudy (anonymous):

Solve 3x^2 = -30x + 24. A) quantity of negative 10 plus or minus square root of 33 all over 2 B) negative 5 plus or minus 2 square root of 33 C) negative 10 plus or minus square root of 33 D) negative 5 plus or minus square root of 33

OpenStudy (anonymous):

You solve this by moving -3x^2 from both sides so you get an equation that looks like this \[0=-3x^2 -30x+24\] And then you use the quadratic formula to get the answer. Do you know the quadratic formula?

OpenStudy (anonymous):

yess

OpenStudy (anonymous):

So what answers do you get when you use quadratic formula?

OpenStudy (anonymous):

\[30\sqrt{612} \div -6\]

OpenStudy (anonymous):

i dont think its right

OpenStudy (anonymous):

You have the 30 and -6 part right, but the square root of 612, I don't think is correct. When you apply the quadratic formula the inside of the square root should look like this \[\sqrt{-30^2-4(-3)(24)}\] When you calculate this through, you should get \[\sqrt{1188}\]

OpenStudy (anonymous):

Oh yeah one more thing I didn't catch, it should be 30 (plus or minus) square root, not 30 times squareroot.

OpenStudy (anonymous):

So it should look like this \[(30(+ or -)\sqrt{1188})/-6\]

OpenStudy (anonymous):

\[5\pm2\sqrt{33}\]

OpenStudy (anonymous):

The correct answer is \[-5\pm \sqrt{33}\] So to me that sounds like D, but I can't look at what the answers look like, but yeah that's it.